# MelCap: A Unified Single-Codebook Neural Codec for High-Fidelity Audio Compression

### v2

@@ 4.3 Audio Reconstruction
Table 2: Objective evaluation metrics for different codecs.
@@
DAC 9 774 4.46 0.67 1.77 0.65 0.87 1.31
Spectral Codec 9 672 4.04 0.90 2.77 0.96 0.64 1.30
Nvidia Codec 9 672 4.05 0.89 2.81 0.96 0.82 1.30
SNAC 4 240 4.35 0.68 1.69 0.68 0.76 1.31
DAC(s) 4 344 4.15 0.80 2.54 0.95 0.78 1.31
WaveTokenizer 1 75 4.20 0.71 2.30 0.76 0.65 1.31
Mel Cap w/o vgg 1 260 4.18 0.76 1.87 0.57 0.52 1.31
Mel Cap w/ vgg 1 260 4.29 0.66 1.90 0.56 0.63 1.31
In the second stage, we use the mel-spectrogram tokens obtained from the first stage as input. After training the second stage, our final results are obtained from the jointly optimized decoder and vocoder. Our evaluation, conducted on the AudioSet test set and detailed in  2. The result shows that Mel Cap with VGG loss achieves competitive perceptual quality (VISQOL 4.29) while using only a single codebook. It also obtains the best fidelity metrics (lowest LSD and Mel Distance), outperforming its non-VGG variant, single-codebook baseline and multi-codebook baselines. This demonstrates that incorporating VGG loss effectively mitigates over-smoothing and enhances spectral reconstruction , which is beneficial for training in second stage.
@@
We conducted a MUSHRA-style listening test to evaluate the perceptual quality of the generated audio. A total of 15 participants were recruited for the experiment. Each participant was presented with 30 randomly selected audio samples drawn from a diverse set of test cases including music, speech, and general sounds. For each trial, participants were asked to rate the audio samples on a continuous quality scale, following the MUSHRA protocol. After the test, we aggregated the ratings across all participants and samples to obtain the final statistical results. The subjective evaluation indicates that Mel Cap achieves perceptual quality comparable to other multi-codec approaches.
@@
Figure 5: Mel-spectrogram comparison of original and reconstructed waveforms produced by different codecs. (a) Ground-truth; (b) DAC; (c) Mel Codec; (d) Nvidia Codec; (e) SNAC; (f) DAC(small); (g) WaveTokenizer; (h) Mel Cap. MelCap accurately reconstructs the high-frequency details.
Table 3: Objective quality on the Music dataset. Best/second-best are marked in bold/underlined.
@@
DAC 9 774 4.20 0.96 2.29 0.41 0.89 1.31
Spectral Codec 9 672 4.31 1.14 2.80 0.55 0.82 1.31
Nvidia Codec 9 672 3.94 1.14 2.83 0.50 0.83 1.30
SNAC 4 240 4.06 1.15 3.09 0.64 0.71 1.30
DAC(s) 4 240 3.78 1.32 3.08 0.64 0.83 1.31
WaveTokenizer 1 75 2.94 1.49 3.74 0.85 0.68 1.32
Mel Cap 1 260 3.97 1.01 2.70 0.40 0.61 1.32
@@
4.4 Downstream Task Evaluation
Unlike speech-only datasets, which can be evaluated using reconstructed waveform quality by ASR models, AudioSet contains multiple sound categories and is designed for audio classification tasks. We further evaluate our codec on the downstream classification task. Specifically, we employ pretrained models from Dinkel et al. (2023) and compute the top-3 predicted labels using the reconstructed waveforms. To assess the codec’s ability to preserve semantic information, we report F1 and mAP scores in Table  4, which measure the accuracy of sound event classification. The performance demonstrates the codec’s effectiveness in retaining discriminative detail beyond perceptual quality, achieving better downstream classification results compared to other codec baselines.
5 Future Work
Currently, open-source high-fidelity audio datasets are scarce, and even 44kHz/48kHz corpora often include upsampled 16kHz content, limiting compression performance. In future work, we plan to collect more high-resolution recordings, enhance our model for better perceptual quality, and develop suitable benchmarks for evaluating non-speech, non-music sounds.
@@ 6 Conclusion
In this paper, we introduce MelCap, a “one-codebook-for-all” neural codec that compresses speech, music, and general sound with a single codebook. We achieved this by converting mel-spectrograms into discrete tokens in a single codebook, and then reconstructing high-quality audio from these tokens using a vocoder. During this process, we encountered over-smoothing issues, which we mitigated through careful loss design. These findings provide new insights for future research in audio codec development. Specifically: First, incorporating perceptual loss into the reconstruction stage helps preserve more structure detail in the codebook. Second, two-stage training stabilizes the overall training process, leading to faster convergence than GAN-based models. Third, using mel-spectrograms as the representation better preserves frequency details during compression.

### v3

@@ 4.3 Audio Reconstruction
Table 3: Objective evaluation metrics for MelTok and baselines on Audioset eval set. We bold the best results in all the models, and bold and underline the best results in single-codebook codec models. (24k: ground truth resampled to 24kHz before measurement)
@@
DAC 9 9.0 774 0.81 2.06 0.50 4.19 3.90 0.03
SNAC 4 2.6 240 0.86 2.07 0.61 4.01 3.80 0.03
Spectral Codec 9 6.9 672 0.99 2.42 0.72 4.02 4.82 0.02
Nvidia Codec 9 6.9 672 1.07 2.45 0.74 3.96 5.39 0.08
Wavtokenizer (audio/music) 1 0.9 75 1.18 2.19 0.45(24k) 3.26 4.88 0.15
Wavtokenizer (unify) 1 0.48 40 1.24 2.37 0.52(24k) 3.26 5.32 0.16
UniCodec 1 0.9 75 1.24 2.11 0.38(24k) 3.18 4.97 0.12
Our Method 1 3.4 260 0.77 2.07 0.44 4.17 2.92 0.12
In the second stage, we use the mel-spectrogram tokens obtained from the first stage as input. After training the second stage, our final results are obtained from the jointly optimized decoder and vocoder. Our evaluation, conducted on the AudioSet test set and detailed in  3. The result shows that Mel Tok achieves competitive perceptual quality (VISQOL 4.17) while using only a single codebook. It also obtains the best fidelity metrics (lowest LSD and MCD Distance), outperforming single-codebook baseline and multi-codebook baselines.
@@
We conducted a MUSHRA-style listening test to evaluate the perceptual quality of the generated audio. A total of 15 participants were recruited for the experiment. Each participant was presented with 30 randomly selected audio samples drawn from a diverse set of test cases including music, speech, and general sounds. For each trial, participants were asked to rate the audio samples on a continuous quality scale, following the MUSHRA protocol. After the test, we aggregated the ratings across all participants and samples to obtain the final statistical results. The subjective evaluation indicates that Mel Tok achieves perceptual quality comparable to other multi-codec approaches.
@@
Table 4: Objective quality on the musdb18-hq test dataset.We bold the best results in all the models, and bold and underline the best results in single-codebook codec models. “Original” refers to results obtained using the ground-truth waveform. (24k: ground truth resampled to 24kHz before measurement)
@@
DAC 0.85 1.71 0.36 4.18 2.04 1.85 0.98
SNAC 1.20 2.53 0.53 3.80 3.08 2.10 0.96
Spectral Codec 1.65 3.17 0.73 3.92 3.12 2.46 0.90
Nvidia Codec 1.77 3.24 0.76 3.84 3.28 2.22 0.92
Wavtokenizer
(audio/music)
2.01 2.20 0.47 (24k) 2.73 3.90 3.31 0.92
Wavtokenizer
2.04 2.39 0.62 (24k) 2.69 4.38 3.06 0.89
UniCodec 1.99 2.12 0.40 (24k) 2.83 3.26 2.34 0.93
Our Method 0.90 2.02 0.40 4.23 2.46 2.14 0.96
@@
Table 6: Downstream music genre classification performance.
[…]
Reference 0.63 0.63 0.58
DAC 0.60 0.57 0.57
SNAC 0.63 0.59 0.57
Nvidia Codec 0.60 0.57 0.57
Spectral Codec 0.65 0.59 0.57
WavTokenizer (audio) 0.27 0.15 0.15
WavTokenizer (unify) 0.24 0.12 0.18
UniCodec 0.26 0.15 0.25
Our Method 0.57 0.50 0.50
@@
4.4 Downstream Audio Understanding Task Evaluation
Unlike speech-only datasets, which can be evaluated using reconstructed waveform quality by ASR models, AudioSet contains multiple sound categories and is designed for audio event classification. We further evaluate our codec on this task. Specifically, we employ pretrained models from Dinkel et al. (2023) and compute the top-3 predicted labels using the reconstructed waveforms. To assess the codec’s ability to preserve semantic information, we report F1 and mAP scores in Table  5, which measure the accuracy of sound event classification. The performance demonstrates the codec’s effectiveness in retaining discriminative detail beyond perceptual quality, achieving better downstream classification results compared to other codec baselines.
Also, we performed music genre classification using a fine-tuned version of the pre-trained model DistilHuBERT  Chang et al. (2022) on the GTZAN dataset  Tzanetakis et al. (2001). As shown in Table 6, our method achieves the best performance among single-codebook codecs, while remaining competitive with multi-codebook codecs such as Spectral Codec. This demonstrates that our approach provides an effective representation for downstream music understanding tasks.
5 Limitations and Future Work
Our experiments reveal that training MelTok is disrupted when using low-resolution data that has been artificially upsampled to a high sampling rate. Currently, open-source high-fidelity audio datasets are scarce, and even 44kHz/48kHz corpora often include upsampled 16kHz content, limiting compression performance. In future work, we plan to collect more high-resolution recordings, enhance our model for better perceptual quality, and develop suitable benchmarks for evaluating non-speech, non-music sounds.
Although our experiments demonstrate that preserving more acoustic detail in the codec is promising for audio understanding tasks, its impact on downstream generation tasks remains unknown. Due to space and computational constraints, we have focused on showcasing MelTok’s reconstruction capabilities and have not yet use tokens obtained from MelTok to train in Audio Language Models (ALM). In future work, we plan to investigate the performance of MelTok-based ALMs on downstream audio tasks.
@@ 6 Conclusion
In this paper, we introduce MelTok, a neural codec that compresses full-bandwidth audio into a single codebook. We achieve this by converting mel-spectrograms into discrete tokens using a 2D tokenizer, and then reconstructing high-quality audio from these tokens with a vocoder. In the process, we investigate the disentangled distributions obtained from 2D tokenization. These findings offer new insights for future research in audio codec development. Specifically: First, incorporating perceptual loss into the reconstruction stage helps preserve more structure detail in the codebook. Second, two-stage training stabilizes the overall training process, leading to faster convergence than GAN-based models. Third, using mel-spectrograms as the representation better preserves frequency details during compression.
@@ A.6 Proof of ISTFT Lipschitz Continuity
A.7 Architecture Details For Tokenization
Table 7: The encoder architecture details of MelTok
Layer Name
Input Shape
(Batch,Channel,Frequency,Time)
Kernel Size, Stride Output Shape
Patcher (8, 1, 96, 96) (8, 1, 48, 48)
PreConv2D (8, 1, 48, 48) (3, 1) (8, 128, 48, 48)
Downsampling 1:
ResnetBlock (8, 128, 48, 48) (3, 1), (1, 1) (8, 256, 48, 48)
ResnetBlock (8, 256, 48, 48) (3, 1), (1, 1) (8, 256, 48, 48)
DownConv2D (8, 256, 48, 48) (2, 2) (8, 256, 24, 24)
AttnBlock (Cond.) (8, 256, 24, 24) (1, 1) (8, 256, 24, 24)
Downsampling 2:
ResnetBlock (8, 256, 24, 24) (3, 1), (1, 1) (8, 512, 24, 24)
ResnetBlock (8, 512, 24, 24) (3, 1), (1, 1) (8, 512, 24, 24)
DownConv2D (8, 512, 24, 24) (2, 2) (8, 512, 12, 12)
AttnBlock (Cond.) (8, 512, 12, 12) (1, 1) (8, 512, 12, 12)
ResnetBlock $\times$ 2 (8, 512, 12, 12) (3, 1), (1, 1) (8, 512, 12, 12)
AttnBlock (Cond.) (8, 512, 12, 12) (1, 1) (8, 512, 12, 12)
Normalizer (8, 512, 12, 12) (8, 512, 12, 12)
PostConv2D (8, 512, 12, 12) (3, 1) (8, 512, 12, 12)
Table 8: 2D Tokenizer vs 1D Tokenizer Architecture in Ablation Study
2D Tokenizer 1D Tokenizer
Input $\times 96 96×96$$\times 96$
Encoder $96\times 96\to 4\times 48\times 48$ (Haar, patch=2) $96\times 96\to 128\times 96$
$4\times 48\times 48\to 128\times 48\times 48$ $128\times 96$$256\times 96$
Downsample: Downsample:
$128\times 48\times 48\to 256\times 24\times 24\to 512\times 12\times 12$ $256\times 96\to 512\times 12$
Pre quant conv $512\times 12\times 12\to 256\times 12\times 12$ $512\times 12\to 256\times 12$
Quant_conv $256\times 12\times 12\to 64\times 12\times 12$ $256\times 12\to 64\times 12$
Quantizer 8192 codes of dim 64 8192 codes of dim 64
One Quantizer 12 Quantizers to keep same bitate
Post quant conv $64\times 12\times 12\to 256\times 12\times 12$ $64\times 12\to 256\times 12$
Decoder Symmetric to Encoder Symmetric to Encoder
Reconstruction $\times 96 96×96$$\times 96$